A 189.3-kg and a 258.9-kg cart approach each other on a horizontal air track. They collide and stick together. After the collision their total kinetic energy is 1142.5 J. What is the speed after the collision

1
Expert's answer
2022-05-20T13:19:45-0400
"E_k=\frac{mv^2}{2}""v=\sqrt{2E_k/m}\\
=\sqrt{2*1142.5/(189.3+258.9)}=2.258\:\rm m/s"